# Trio-ViT: Post-Training Quantization and Acceleration for Softmax-Free Efficient Vision Transformer

### v1

@@ VI-A Experimental Setup
Dataset, Baselines, and Metrics. We validate our Trio-ViT’s post-training quantization algorithm on the ImageNet dataset [37]. Specifically, we randomly sample images from the training set as calibration data and then test on the validation set. ❶ To verify the effectiveness of our quantization engine, we consider seven baselines: MinMax, EMA [38], Percentile [39], OMSE [40], Bit-Split [41], EasyQuant [42], and the SOTA method FQ-ViT [9], for standard ViTs[4]/DeiTs[5], and compare with them in terms of accuracy (top-1 accuracy by default). ❷ To validate our dedicated accelerator, we consider nine baselines: (i) full-precision ViTs executed on the widely-used Edge GPU (NVIDIA Tegra X2); -bit ViTs quantized following (ii) FasterTransformer [43], (iii) I-BERT [44], and (iv) I-ViT [11], and accelerated on the Turing Tensor Core of GPU (NVIDIA 2080Ti), which supports efficient integer arithmetic via TVM; and full-precision EfficientViT executed on the (v) widely-used Edge CPU (Qualcomm Snapdragon 8Gen1 CPU) and Edge GPUs, including (vi) NVIDIA Jetson Nano and (vii) NVIDIA Jetson Orin; as well as two SOTA ViT accelerators, including (viii) the dedicated accelerator Auto-ViT-Acc [16] for standard ViTs and (ix) ViA [45] tailored for Swin Transformer [22], one of efficient ViTs. We compare them in terms of throughput, energy efficiency, frame rate, and DSP efficiency.
Accelerator Setup. Characteristics: The parallelism of computing engines in our accelerator (as depicted in Fig. 8) is configured to as . Thus, there are a total of multipliers in our accelerator, each can execute an -bit multiplication. To improve DSP utilization, we adopt the SOTA DSP packing strategy [46] to accommodate two -bit multiplications within each DSP, similar to Auto-ViT-Acc [16] for fair comparisons. Evaluation: We implement our accelerator with Verilog, synthesize through Vivado Design Suite, and evaluate on Xilinx ZCU102 FPGA at 200-MHz frequency. Table III lists our resource consumptions. Furthermore, we develop a cycle-accurate simulator for our accelerator to obtain fast and reliable estimations and verify them against the RTL implementation to ensure correctness.
@@ VI-B Evaluation of Trio-ViT’s Post-Training Quantization
Results and Analysis. From Table IV, we can draw four conclusions. (i) The SOTA post-training quantization (PTQ) method FQ-ViT [9], which develops dedicated quantization schemes to fully quantize all operations in standard ViTs (including Softmax) to enhance hardware efficiency, suffers from accuracy compared to the full-precision models. Besides, it also yields an average accuracy compared to the base PTQ, where Softmax and other non-linear operations are not quantized, thus incurring non-negligible hardware costs. This demonstrates that the hardware-unfriendly non-linear operations are sensitive to quantization, hindering both the achievable hardware efficiency and quantization accuracy of standard ViTs. (ii) To address this limitation, the SOTA efficient ViT dubbed EfficientViT [12] has been proposed, which features Softmax-free linear attention (LinAttn) and can achieve much higher accuracy with even fewer parameters and computational costs. This underscores EfficientViT’s superiority, highlighting the need for quantization to facilitate its real-world applications. (iii) However, due to the distinct distributions of activation in MBConvs and MSAs, as introduced in Sec. IV-A, the vanilla PTQ method fails to quantize EfficientViT and even yields a Not-a-Number (NaN) issue. (iv) To solve this issue, we propose our dedicated PTQ engine, which can effectively quantize EfficientViTs with merely an average accuracy when compared with the full precision counterparts, demonstrating our effectiveness.
@@
Effectiveness of Our Dedicated Quantization Engine. As shown in Table V, we can see that: As for the quantization within MBConvs, (i) due to the inter-channel variations in DW’s inputs and inter-channel asymmetries in PW2’s inputs, as introduced in Sec. IV-A1, vanilla uniform quantization fails to quantize MBConvs in EfficientViT. (ii) By incorporating our channel-wise migration and filter-wise shifting, which are proposed to solve the above two issues, respectively, and are both indispensable, we can effectively quantize MBConvs with only an average accuracy. On top of this, regarding the quantization of lightweight MSA, (iii) owing to the extreme quantization sensitivity of smaller values in divisors, as illustrated in Sec. IV-A2, the vanilla -bit uniform quantization yields a NaN issue. (iv) Thus, we advocate adopting log2 quantization for divisors, which assigns more bins to smaller values and is inherently compatible with the algorithmic property of divisors, thus allowing for the effective quantization of divisors with a mere -bit.
@@
Effectiveness of Proposed Channel-Wise Migration. Considering the unique algorithmic property of DWConvs, where each channel of weights serves as an independent filter to process each input channel, channel-wise quantization serves as a straightforward solution to solve the inter-channel variations in DW’s input, as explained in Sec. IV-B. However, it will greatly increase the number of scaling factors, challenging quantization optimization via LSQ [32] and limiting achievable accuracy. Thus, we advocate adopting channel-wise migration on top of layer-wise quantization. As shown in Table VI, by doing this, we can offer an average accuracy, demonstrating our superiority in solving the inter-channel variations of DW’s input while maintaining optimization efficiency.
@@ VI-C Evaluation of Trio-ViT’s Dedicated Accelerator
Comparisons with SOTA Baselines on GPUs/CPUs. We follow [47, 26] to scale up the hardware resources of our accelerator to have comparable peak throughput with the general computing platform (i.e., NVIDIA 2080Ti GPU) to enable fair comparisons with GPUs/CPUs. As shown in Fig. 10 (where the y-axis is plotted on a logarithmic scale for better illustration), we can achieve much better hardware efficiency compared to SOTA baselines on GPUs/CPUs, validating our effectiveness. Specifically, (i) when compared with the full-precision DeiTs [5] on Edge GPU (Tegra X2), we can achieve and throughput and energy efficiency, respectively. (ii) As for comparisons with 8-bit DeiTs quantized by FasterTransformer [43], I-BERT [44] and I-ViT [11], and executed on the Turning Tensor Core of the NVIDIA 2080Ti GPU, we can offer and throughput and energy efficiency, respectively. Furthermore, (iii) regarding the comparison with full precision EfficientViT [12] on edge CPU, we can gain up to and throughput and energy efficiency, respectively. (iv) When compared with EfficientViT on edge GPUs (NVIDIA Jetson Nano and Jetson Orin), we can gain and in terms of throughput and energy efficiency.
| Accelerator | Via [45] | Auto-ViT-Acc [16] | Ours | |||
| Device | Xilinx | Xilinx | Xilinx | |||
| Alveo U50 | ZCU102 | ZCU102 | ||||
| Frequency (MHz) | 300 | 150 | 200 | |||
| Format | FP16 | FIX8 | FIX8 | |||
| Model | Swin-T | Deit-Small | Deit-Base | EfficientViT | EfficientViT | EfficientViT |
| -R224 | -R224 | -R224 | -B2-R288 | -B1-R224 | -B2-R256 | |
| DSP Used | 2420 | 1936 | 2066 | 1024 | 1024 | 1024 |
| Latency (ms) | 14.5 | 12.8 | 38.6 | 1.12 | 2.02 | 2.64 |
| FPS | 68.8 | 78.1 | 25.9 | 895 | 494 | 379 |
| Throughput | 310 | 711 | 900 | 769 | 791 | 796 |
| DSP Efficiency | 0.13 | 0.37 | 0.44 | 0.75 | 0.77 | 0.78 |
| (GOPS/DSP) | ||||||
| Energy Efficiency | 7.92 | 84.1 | 95.7 | 105 | 108 | 109 |
| (GOPS/W) | ||||||
| Accuracy (%) | 81.3 | 79.69 | 81.93 | 79.58 | 80.97 | 81.62 |
@@
Comparisons with SOTA ViT Accelerators. From Table VII, we can see that: (i) due to the superiority of our quantization engine and the promising hardware efficiency of EfficientViT compared to standard DeiT, we gain the lowest latency and highest frame rate (FPS) under comparable accuracy. Particularly, we can achieve up to FPS compared to Via [45], a dedicated accelerator for an efficient ViT dubbed Swin-Transformer-Tiny (Swin-T) [22], as well as FPS over the dedicated ViT/DeiT accelerator Auto-ViT-Acc [16]. (ii) Furthermore, owing to the hybrid design of our dedicated accelerator, which is proposed to effectively support various operators in the Convolution-Transformer hybrid architecture of EfficientViT with enhanced hardware utilization, we can achieve the highest hardware utilization efficiency. For example, we can offer up to and DSP efficiency when compared with Via and Auto-ViT-Acc, respectively. (iii) Additionally, thanks to our developed pipeline architecture aiming to facilitate both inter- and intra-layer fusion, we gain the best energy efficiency, i.e., up to and compared to Via and Auto-ViT-Acc, respectively.
@@ VII Conclusion
In this paper, we have proposed, developed, and validated Trio-ViT, the first post-training quantization and acceleration framework dedicated to the state-of-the-art (SOTA) efficient Vision Transformer (ViT), dubbed EfficientViT. Specifically, at the algorithm level, we propose a tailored post-training quantization engine that incorporates several innovative quantization schemes to effectively quantize EfficientViT with enhanced quantization accuracy. At the hardware level, we develop a dedicated accelerator integrating a hybrid design and a pipeline architecture, aiming to boost hardware efficiency. Extensive experimental results consistently prove our effectiveness. Particularly, we can gain up to and frame rate with comparable accuracy over state-of-the-art ViT accelerators.

### v3

@@ VI-A Experimental Setup
Dataset, Baselines, and Metrics. We validate our Trio-ViT’s post-training quantization algorithm on the ImageNet dataset [38] and implement it on the NVIDIA GeForce RTX 3090 GPUs, each with GB of memory. Specifically, we randomly sample images from the training set as calibration data and then test on the validation set. ❶ To verify the effectiveness of our quantization engine, we consider seven baselines: MinMax, EMA [39], Percentile [40], OMSE [41], Bit-Split [42], EasyQuant [43], and the SOTA method FQ-ViT [10], for standard ViTs[4]/DeiTs[5], and compare with them in terms of top-1 accuracy. ❷ To validate our dedicated accelerator, we consider ten baselines: (i) full-precision ViTs executed on the widely-used Edge GPU (NVIDIA Tegra X2); -bit ViTs quantized following (ii) FasterTransformer [44], (iii) I-BERT [45], and (iv) I-ViT [12], and accelerated on the Turing Tensor Core of NVIDIA 2080Ti GPU, which supports efficient integer arithmetic via TVM; and full-precision EfficientViT executed on the (v) widely-used Edge CPU (Qualcomm Snapdragon 8Gen1) and Edge GPUs, including (vi) NVIDIA Jetson Nano and (vii) NVIDIA Jetson Orin; as well as three SOTA ViT accelerators, including (viii) Auto-ViT-Acc [17] and (ix) Huang et al. [46] tailored for standard ViTs and (x) ViA [47] dedicated to Swin Transformer [23] (one of efficient ViTs). We compare them in terms of throughput, energy efficiency, frame rate (FPS), and DSP efficiency.
Accelerator Setup. Characteristics: The parallelism of computing engines in our accelerator (as depicted in Fig. 8) is configured to as . Thus, there are a total of multipliers in our accelerator, each can execute an -bit multiplication. To improve DSP utilization, we adopt the SOTA DSP packing strategy [48] to accommodate two -bit multiplications within each DSP, similar to Auto-ViT-Acc [17] for fair comparisons. Evaluation: We implement our accelerator with Verilog, synthesize through Vivado Design Suite, and evaluate on Xilinx ZCU102 FPGA at 200-MHz frequency. Table III lists our resource consumptions. Furthermore, we follow [28, 9] to develop a cycle-accurate simulator for our accelerator to obtain fast and reliable estimations and verify them against the RTL implementation to ensure correctness.
@@ VI-B Evaluation of Trio-ViT’s Post-Training Quantization
Results and Analysis. From Table IV, we can draw four conclusions. (i) The SOTA post-training quantization (PTQ) method FQ-ViT [10], which develops dedicated quantization schemes to fully quantize all operations in standard ViTs (including Softmax) to enhance hardware efficiency, suffers from accuracy compared to the full-precision models. Besides, it also yields an average accuracy compared to the base PTQ, where Softmax and other non-linear operations are not quantized and thus incur non-negligible hardware costs. This demonstrates that the hardware-unfriendly non-linear operations are sensitive to quantization, hindering both the achievable hardware efficiency and quantization accuracy of standard ViTs. (ii) To address this limitation, the SOTA efficient ViT dubbed EfficientViT [13] has been proposed, which features Softmax-free linear attention (LinAttn) and can achieve much higher accuracy with even fewer parameters and computational costs. This underscores EfficientViT’s superiority, highlighting the need for quantization to facilitate its real-world applications. (iii) However, due to the distinct distributions of activation in MBConvs and MSAs, as introduced in Sec. IV-A, the vanilla PTQ method fails to quantize EfficientViT and even yields a Not-a-Number (NaN) issue. (iv) To solve this issue, we propose our dedicated PTQ engine, which can effectively quantize EfficientViTs with merely an average accuracy when compared with the full precision counterparts, demonstrating our effectiveness.
@@
Effectiveness of Our Dedicated Quantization Engine. As shown in Table V, we can see that: As for the quantization within MBConvs, (i) due to the inter-channel variations in DW’s inputs and inter-channel asymmetries in PW2’s inputs, as introduced in Sec. IV-A1, vanilla uniform quantization fails to quantize MBConvs in EfficientViT. (ii) By incorporating our channel-wise migration and filter-wise shifting, which are proposed to solve the above two issues, respectively, we can effectively quantize MBConvs with only an average accuracy. On top of this, regarding the quantization of lightweight MSA, (iii) owing to the extreme quantization sensitivity of smaller values in divisors, as illustrated in Sec. IV-A2, the vanilla -bit uniform quantization yields a NaN issue. (iv) Thus, we advocate adopting log2 quantization for divisors, which assigns more bins to smaller values and is inherently compatible with the algorithmic property of divisors, thus allowing for quantizing divisors with a mere -bit.
@@
Effectiveness of Proposed Channel-Wise Migration. Considering the unique algorithmic property of DWConvs, where each channel of weights serves as an independent filter to process each input channel, channel-wise quantization serves as a straightforward solution to solve the inter-channel variations in DW’s input, as explained in Sec. IV-B. However, it will greatly increase the number of scaling factors, challenging quantization optimization via LSQ [33] and limiting achievable accuracy. Thus, we advocate adopting channel-wise migration on top of layer-wise quantization. As validated in Table VI, by doing this, we offer an average accuracy, demonstrating our superiority in solving the inter-channel variations of DW’s input while maintaining optimization efficiency.
| Methods | MBConv Quant | MSA Quant | EfficientViT -B0-R1024 | |||
| Vanilla | CW Migration | FW Shifting | Uniform (8) | Log2 (4) | ||
| Full Precision | – | – | – | – | – | 75.65 |
| Base PTQ | ✓ | ✓ | N/A | |||
| Ablation Study in MBConv | ✓ | – | – | 36.89 | ||
| ✓ | – | – | 72.51 | |||
| ✓ | – | – | 29.02 | |||
| ✓ | ✓ | – | – | 75.22 | ||
| Ablation Study in MSA | ✓ | ✓ | ✓ | 1.12 | ||
| ✓ | ✓ | ✓ | 74.83 | |||
Generalization on Semantic Segmentation. To assess the generalization capability of our proposed quantization method, we apply it to EfficientViT-B0-R1024 [13] and evaluate its performance on the semantic segmentation task, using the Cityscapes [49] as the dataset and mean Intersection over Union (mIoU) as the evaluation metric. As listed in Table VII, we can see that: For the quantization within MBConvs, vanilla uniform quantization fails due to the inter-channel variations in DW’s inputs and inter-channel asymmetries in PW2’s inputs. By combining our channel-wise (CW) migration and filter-wise (FW) shifting, we achieve effective quantization of MBConvs with only a reduction in mIoU. Note that, due to the interdependence between layers, while filter-wise shifting alone is inferior to the vanilla approach, it can boost performance when incorporated with our channel-wise migration, proving its effectiveness and necessity. Furthermore, regarding the quantization of lightweight MSA, owing to the extreme quantization sensitivity of smaller values in divisors, the vanilla -bit uniform quantization yields a catastrophic performance drop. In contrast, our proposed log2 quantization enables the effective quantization of divisors with a mere -bit. This set of results demonstrates the generalization ability and robustness of our quantization method.
@@ VI-C Evaluation of Trio-ViT’s Dedicated Accelerator
Comparisons with GPUs/CPU. We follow [50, 27] to scale up the hardware resources of our accelerator to have comparable peak throughput with the general computing platform (i.e., NVIDIA 2080Ti GPU) to enable fair comparisons with GPUs/CPU. As shown in Fig. 10 (where the y-axis is plotted on a logarithmic scale for better illustration), we can achieve much better hardware efficiency compared to SOTA baselines on GPUs/CPU, validating our effectiveness. Specifically, (i) when compared with the full-precision DeiTs [5] on Edge GPU (Tegra X2), we can achieve and throughput and energy efficiency, respectively. (ii) For comparisons with 8-bit DeiTs quantized by FasterTransformer [44], I-BERT [45] and I-ViT [12], and executed on the Turning Tensor Core of NVIDIA 2080Ti GPU, we can offer and throughput and energy efficiency, respectively. Furthermore, (iii) regarding the comparison with full precision EfficientViT [13] on edge CPU, we can gain up to and throughput and energy efficiency, respectively. (iv) When compared with EfficientViT on edge GPUs (NVIDIA Jetson Nano and Jetson Orin), we can gain and in terms of throughput and energy efficiency.
| Accelerator | Via [47] | Huang et al. [46] | Auto-ViT-Acc [17] | Ours | ||||||||||||||||||||
| Device |
| Frequency (MHz) | 300 | 300 | 150 | 200 | ||||||||||||||||||||
| Format | FP16 | INT8 | INT8 | INT8 | ||||||||||||||||||||
| Model |
| DSP Used | 2420 | 1268 | 1268 | 1936 | 2066 | 1024 | 1024 | 1024 | ||||||||||||||||
| Latency (ms) | 14.5 | 4.1 | 11.2 | 12.8 | 38.6 | 2.24 | 4.05 | 5.28 | ||||||||||||||||
| Frame Rate (FPS) | 68.8 | 245 | 89.3 | 78.1 | 25.9 | 447 | 247 | 190 | ||||||||||||||||
310 | 616 | 763 | 711 | 900 | 769 | 791 | 796 | ||||||||||||||||
0.13 | 0.49 | 0.60 | 0.37 | 0.44 | 0.75 | 0.77 | 0.78 | ||||||||||||||||
7.92 | – | 25.8 | 84.1 | 95.7 | 105 | 108 | 109 | ||||||||||||||||
| Accuracy (%) | 81.3 | 74.39 | – | 79.69 | 81.93 | 79.58 | 80.97 | 81.62 | ||||||||||||||||
@@
Comparisons with SOTA ViT Accelerators. From Table VIII, we can see that: (i) Due to the superiority of our quantization engine and the promising hardware efficiency of EfficientViT compared to standard ViTs/DeiTs, we gain the lowest latency and highest FPS under comparable accuracy. Particularly, we achieve up to FPS compared to Via [47], a dedicated accelerator for an efficient ViT dubbed Swin-Transformer-Tiny (Swin-T) [23], as well as and FPS over dedicated ViT/DeiT accelerators Huang et al. [46] and Auto-ViT-Acc [17], respectively. (ii) Furthermore, owing to the hybrid design of our dedicated accelerator, which is proposed to effectively support various operators in the Convolution-Transformer hybrid architecture of EfficientViT with enhanced hardware utilization, we can achieve the highest hardware utilization efficiency. For example, we can offer up to , , DSP efficiency when compared with Via, Huang et al., and Auto-ViT-Acc, respectively. (iii) Additionally, thanks to our developed pipeline architecture aiming to facilitate both inter- and intra-layer fusion, we gain the best energy efficiency, i.e., up to , , and compared to them.
@@ VII Conclusion
In this paper, we have proposed, developed, and validated Trio-ViT, the first post-training quantization and acceleration framework dedicated to the state-of-the-art (SOTA) efficient Vision Transformer (ViT), dubbed EfficientViT. Specifically, at the algorithm level, we propose a tailored post-training quantization engine that incorporates several innovative quantization schemes to effectively quantize EfficientViT with enhanced quantization accuracy. At the hardware level, we develop a dedicated accelerator integrating a hybrid design and a pipeline architecture to boost hardware efficiency. Extensive experimental results consistently prove our effectiveness. Particularly, we gain up to , , and FPS with comparable accuracy over SOTA ViT accelerators.